OpenStudy (anonymous):

What is the perimeter of a triangle with vertices located at (-1, 4), (2, 7), and (1, 5), rounded to the nearest hundredth? A. 7.63 units B. 7.89 units C. 8.71 units D. 9.24 units

OpenStudy (anonymous):

i think C

OpenStudy (amilapsn):

\[\sqrt{(-1-2)^2+(4-7)^2}+\sqrt{(-1-1)^2+(4-5)^2}+\sqrt{(2-1)^2+(7-5)^2}\]

OpenStudy (anonymous):

wow

OpenStudy (amilapsn):

Do the calculation yourself... Try to remember the formula...

OpenStudy (anonymous):

it is C